Juan invested part of his money at 10% and the rest at 12%. The income from both investments was $2700. If the amount invested at 10% is $5000 less than twice the amount invested at 12%, find the amount that he has invested at each rate.
Answer by Boreal(15235)

x= amount invested at 12%
2x-5000= amount invested at 10%
Therefore .12x+.10(2x-5000)=2700
.12x+..20x-500=2700
.32x=3200
x=$10000 @ 12% =$1200
2x-5000=$15000@10%=$1500
Total is $2700.